OpenStudy (anonymous):

Kara has 21 collectible buttons to sell. She sells the small buttons for $3 and the large buttons for $4, and earns a total of $68. If the small buttons are represented by x and the large buttons are represented by y, which of the following systems of equations can be used to calculate the number of large and small buttons sold?

OpenStudy (anonymous):

need help!

OpenStudy (anonymous):

16 small and 5 large.

OpenStudy (anonymous):

Set up a proportion table

OpenStudy (anonymous):

x + y = 21, 3x + 4y = 68 x + y = 68, 3x + 4y = 21 x + y = 21, 4x + 3y = 68 x + y = 68, 4x + 3y = 21 those are my options

OpenStudy (anonymous):

16 times 3 equals 48 and 5 times 4 equals 20.

OpenStudy (anonymous):

@Mr.Confused can u help me out when ur done ?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

wait im confused

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

thats not one of the options

OpenStudy (anonymous):

What are the options?